# Hot Rocks Survey III: A deep eclipse for LHS 1140c and a new Gaussian process method to account for correlated noise in individual pixels

### v1

@@ 3.1 Simulating the target signal
The simulations were based on the first eclipse of LHS 1140c, which consisted of 1,262 integrations and covered 233 minutes in total (for more details see Section 4), although we assume the first 30 minutes have been clipped. To minimise computational cost, we simulated eight times fewer integrations. This perimitted each pixel-fitting analysis to run in 40 minutes on a CPU cluster node, and we could run eight of these analyses in parallel. As we were simulating observations with eight times the integration time, we reduced the amplitude of Poisson noise by a factor of to maintain similar constraints on the eclipse depth.
@@
The eclipse model was generated using the package jaxoplanet (Hattori et al. 2024) which implements the equations of Mandel & Agol (2002) in JAX (Bradbury et al. 2018). The eclipse model was fixed for all simulations with the mean values of , , , from literature values of LHS 1140c (as listed in Table 11). The central eclipse time was fixed to the midpoint of the time-series and the eccentricity was fixed to zero. For all fits of the simulated data, the eclipse depth was the only eclipse model parameter being fit for with every other eclipse parameter held fixed to the true values.
@@ 3.2 Simulating white noise and systematics
White noise was generated with the amplitude for each pixel determined by taking the standard deviation of each pixel time-series in the first LHS 1140c eclipse. This noise was then convolved with the IPC convolution kernel from Equation 11 using values of , and , before then being added to the simulated signal. Background noise was also introduced to simulate residual background scatter not removed by a background subtraction. Each integration had a constant value added to all pixels, drawn from DN/s, which given the flux in the target signal works out to adding scatter with amplitude of ppm for a 5px radius circular aperture. We also added scatter which was constant along each row and column but independent between each row and column. This was added by drawing from DN/s for the rows and DN/s for the columns. Our analysis of the real observations did not constrain the presence of these row/column systematics but it was a possibility we wanted to include which both aperture photometry and the pixel-fitting method should be able to account for.
@@
We simulated the flux-conserved systematics by taking a random draw from a GP with an exponential kernel and a length scale of 35 minutes and adding it to one pixel and subtracting it from a neighbouring pixel. This was performed for each pair of neighbouring pixels (i.e. 8 times per pixel) and the amplitude of the draw was scaled to be the geometric mean of the flux on each of the pair of pixels multiplied by the constant ppm. By adding and subtracting the same signal, it means that these systematics are flux-conserving with the only non-flux-conserving effects due to the finite area of the aperture (i.e. there is ”flux movement” across the fixed aperture boundary), although the amplitude of this effect was small. For a pixel surrounded by eight pixels with the same flux values, the amplitude of systematics in this pixel would work out to ppm. Note that was meant to be set to 700 ppm - consistent with the values recovered in the observations - but the addition and subtraction was accidentally performed twice for each pair of pixels, effectively increasing the amplitude of this effect by. We generated 200 of these simulations and we refer to them as the Flux-Conserved Systematics (FCS) Scenario.
The other two sets of simulations included additional systematics which were not flux-conserved. The second set of simulations - called the Independent Pixel Systematics (IPS) Scenario - took each simulation from the FCS Scenario and injected a ppm height scale systematic into a single pixel time-series. The pixel was randomly chosen for each simulation to be one of the nine brightest pixels at the centre of the PSF. As these pixels tend to make up about of total flux, this results in a ppm systematic in the aperture extracted light curve. While aperture extraction with this pixel masked out may be a good approach, the flux-conserved systematics make it difficult to distinguish these additional systematics from flux-conserved systematics. This can be seen in Figure 3, which features a plot of the pixel light curves centred on the PSF for the first simulated dataset in the FCS Scenario and the IPS Scenario.
@@
The third set of simulations added a common systematic shape to all pixels at an amplitude equal to ppm multiplied by the flux on each pixel. These will be referred to as the Common Systematics (CS) Scenario. This could represent real astrophysical systematics such as a flare or instrumental systematics which affect the throughput of the observation. The height scale was chosen as it resulted in a similar amplitude of systematics in an aperture extracted light curve as the IPS Scenario, but with the systematics spread out over all pixels instead of originating entirely from a single pixel.
@@
For each scenario, the synthetic data was aperture extracted with a radius of 5 pixels centred on the best-fit centroid for each integration - as performed for the real observations. This radius minimised scatter in the aperture extracted light curves compared to radii of 4.5 or 5.5 pixels, matching the real observations. See Figure 4 for an example of a simulated dataset from the FCS and IPS Scenarios after aperture extraction. The pixel-fitting method only included pixels with centres within 5 pixels from the PSF centre, resulting in 81 pixels being included, which is similar to the 78.5px2 area for a circular aperture with a radius of 5 pixels.
@@ 3.3 Different fitting approaches tested
We used MCMC to recover the eclipse depth with each method, modelling the posterior as Gaussian and using the mean and standard deviation of the chains. For each method we ran four chains of No U-Turn Sampling (NUTS; Hoffman & Gelman 2014) as implemented in PyMC (Salvatier et al. 2016) - with each chain having 1000 tuning steps and 1000 draws. We also used slice sampling (Neal 2003) in combination with NUTS, by performing blocked Gibbs sampling (Jensen et al. 1995) on some parameters close to their prior bounds (e.g. , , , , , ). This was performed because it was found that running NUTS alone could sometimes result in chains getting stuck on these parameters. The Gelman-Rubin statistic (Gelman & Rubin 1992) was used to measure convergence of the chains. A slightly higher than normal threshold of was used to determine convergence for each simulation in order to reduce computational costs arising from the large number of MCMCs performed.
For each set of simulations, we tested aperture extraction with and without a GP. These MCMCs fit for the eclipse depth , the flux out-of-(secondary) transit , the white noise amplitude and in the case of the GP fits also the height scale and length scale of the GP. We also tested using a GP with fixed to the true simulated value of 35 minutes, as this may be a fairer comparison to the pixel-fitting method - which gains extra information about the time length scale as it fits for flux-conserved systematics with the same time length scale.
For all pixel-fitting approaches, we fit for the eclipse depth , baseline flux for each pixel (for 81 pixels), white noise amplitude for each pixel , change in centroid position at each integration and (for 137 integrations), IPC parameters , and , background scatter parameters , and , height scale and time length scale of flux-conserving systematics and . This describes the ”only flux-conserving systematics” approach (abbreviated ”only FCS”) and includes 445 free parameters which were all fit for. We also included a height scale for systematics with a common shape in all pixels for the common systematics approach and height scales for independent systematics in each pixel for the independent systematics approach. Finally, as this independent systematics approach turned out to give quite conservative predictions, we also reran each of these approaches where only the pixels recovered to have large height scales (specifically where the percentile value was greater than a threshold of ppm) were fit with independent height scales and the remaining pixels all shared a joint height scale . Note these pixel systematics were still treated as independent but with a shared amplitude (i.e. they could still have completely different shapes). This approach is listed in tables as shared independent systematics. We also included methods which could account for both common and independent pixel systematics as these were robust to a range of systematics.
@@ 3.4 Comparing the performance of each method
We measure the precision of each method using the Root Mean-Squared Error (RMSE) of the mean eclipse depth for each simulation. The smaller the RMSE, the more precise the eclipse depth measurement is. If the method is robust, then the average standard deviation of the eclipse depth measurement should approximately match the RMSE. If the standard deviation is too big then our results are overly conservative and if it is too small then the results are unreliable. This can be summarised with the reduced chi-squared statistic , which should equal unity when the uncertainties are correctly estimated, less than one for uncertainties which are too large/conservative and greater than one for uncertainties which are too small/unreliable333This assumes the posteriors for the eclipse depths are Gaussian. For each scenario (consisting of 200 simulations), it is calculated as follows:
Where ppm is the true simulated eclipse depth and , are the mean and standard deviation of the recovered eclipse depth for the simulation.
@@ 3.5 Flux-Conserved Systematics (FCS) Scenario results
This scenario only contained flux-conserved systematics, therefore aperture extraction should result in very clean light curves and a GP should not be required to robustly analyse these observations. However, for real observations we can never be sure that the light curves are completely free from systematics and so we are interested in how well each method performs in this scenario. The methods which marginalise over systematics even though they are not present will naturally tend to give more conservative constraints on the eclipse depths, this is true both for aperture extraction with a GP and for various pixel-fitting approaches. However, this will be balanced out by their more reliable performance when systematics are present.
The results in Table 1 demonstrate this very clearly. The most precise methods (i.e. smallest RMSE) are aperture extraction without a GP and the pixel-fitting approach which only accounts for flux-conserved systematics. They both have values close to unity across these simulations, showing their uncertainty estimates are reliable. Aperture extraction with a GP has a slightly worse precision and larger uncertainties. The significantly smaller value of shows that this method is quite conservative. Pixel-fitting only accounting for common systematics performs similarly.
However, we see that the other pixel-fitting approaches which can account for independent systematics in each pixel have worse accuracy and even larger uncertainty estimates. This is similar to aperture extraction with a GP and in general we do expect a more flexible systematics model to result in larger errorbars as it marginalises over a wider space of models (Gibson 2014). The pixel-fitting method is particularly flexible as it can fit for potential systematics in every pixel. Despite the vast majority of these pixel systematics height scales being consistent with the minimum prior bound, they still marginalise over a range of amplitudes which are consistent with the data. See Figure 5 for the posterior of each independent pixel systematics height scale for the first simulation in the FCS and IPS Scenarios.
We see that sharing the amplitude of the smallest independent pixel systematics height scales does reduce how conservative the method is, with smaller average uncertainties and a value closer to one. Accounting for common systematics as well as independent pixel systematics performs the most conservatively, however this method is the most flexible and can be robustly applied to all three sets of simulations generated.
@@ 3.6 Independent Pixel Systematics (IPS) Scenario results
For aperture extraction with a GP, the height scale of systematics is not constrained to be greater than the minimum prior bound in many of these simulations. This means that varying the location of this prior bound will increase or decrease the weighting of the area of the posterior where systematics are negligible. For example, we get if we increase the minimum bound to and if we instead lower the minimum bound to , we get . Increasing the minimum prior bound on the height scale can be interpreted as a more conservative analysis where the presence of systematics is more likely, and this would help lower the here. However, setting also lowers the value for the first set of simulations to just , which was already overly conservative at . Overall for our normal prior bound of , the from combining the FCS and IPS Scenarios is when fitting for and when fixing it, so the method performs well over an even mix of clean and systematics-contaminated datasets.
The pixel-fitting methods which account for common systematics on top of accounting for independent pixel systematics have slightly lower precision and larger uncertainties. However, they are accounting for additional systematics that could potentially exist in real data - such as from variations in throughput/detector settling - so this is a more conservative approach which might make sense when we do not know the origin of systematics in the data.
@@ 3.7 Common Systematics (CS) Scenario results
Finally, the simulations in the Common Systematics Scenario differ from the IPS Scenario in that the systematics were injected into all pixels with a common shape instead of into a single pixel. The amplitude of these systematics was scaled by the flux in each pixel, so this could represent systematics inherent from the signal, such as stellar granulation or flaring. It could also arise from factors affecting the throughput which vary over time. The amplitude of this was such that aperture extraction with a GP performs similarly to the IPS Scenario but with a different systematics origin. In the IPS Scenario, the pixel-fitting approaches could outperform aperture extraction by weighting towards the cleaner pixels, however this will not work when all the pixels suffer from a similar proportion of systematics. Pixel-fitting should therefore perform similarly to aperture extraction with a GP.
@@ 3.8 Summary of the simulations
The results demonstrate that pixel-fitting can outperform aperture extraction when systematics are large in a particular pixel, but the method is overly conservative when only flux-conserving systematics are present. It also matches the performance of aperture extraction with a GP when systematics have a common shape across all pixels. See Figure 7 for a summary of the results in Tables 1, 2 and 3 for aperture extraction with and without a GP as well as for the general pixel-fitting approach of accounting for common systematics across all pixels and systematics in individual pixels. We can see that compared with aperture extraction with a GP, pixel-fitting performs very similarly when systematics are common between all pixels but much better when there are systematics present in specific pixels - e.g. after a cosmic ray hit (as shown in Figure 1).
@@ 4 Observations
The three eclipses of LHS 1140c were observed on November 2023, July 2024 and July 2024. Each eclipse used the SUB256 subarray, read 36 groups for each integration and had 1,262 integrations. The integrations had an 11.1s cadence, with each observation lasting 233 minutes. This followed a strategy of leaving 60 minutes of flexible start time plus 30 minutes of settling time followed by observing for two eclipse durations, with one eclipse duration ( minutes) before and after expected mid-eclipse for a circular orbit. There was also an additional 9 minutes of padding to help catch slightly more eccentric orbits.
@@ 4.1 Primary data reduction
Our reduction began by downloading the uncalibrated ‘uncal’ FITS files for our three eclipses from the Mikulski Archive for Space Telescopes (MAST), which had already been pre-processed by version 2024_1a of the Science Data Processing (SDP) system. We ran the JWST calibration pipeline version 1.15.1 with the default settings for steps 1 and 2 except we changed the jump detection threshold to 7, excluded both the first and last groups from each integration when reading up-the-ramp (the default is to only exclude the last group) and skipped the photom step in stage 2 to keep the flux units in data numbers per second (DN/s). We excluded the first and last groups to avoid a known issue where these groups can include difficult to correct reset effects (Morrison et al. 2023). The jump detection threshold was decided upon by calculating the average amplitude of white noise recovered across all three eclipse fits when only white noise and an exponential ramp + linear model was fit. It was found that 7 minimised this out of integer thresholds from .
One issue we encountered was that running the pipeline separately on each uncal data segment resulted in offsets in the flux time-series between segments, despite these segments not representing any actual change in the observation but instead are simply a way of reducing the total filesize of any given file. To remedy this, for all our reductions we first combined the uncal segments together into a single FITS file and then ran the JWST pipeline on this combined uncal file, which avoided these offsets.

### v2

@@ 3.1 Simulating the target signal
The simulations were based on the first eclipse of LHS 1140c, which consisted of 1,262 integrations and covered 233 minutes in total (for more details see Section 4), although we assume the first 30 minutes have been clipped. To minimise the computational cost, we simulated eight times fewer integrations. This permitted each pixel-fitting analysis to run in 40 minutes on a CPU cluster node, and we could run eight of these analyses in parallel. As we were simulating observations with eight times the integration time, we reduced the amplitude of Poisson noise by a factor of to maintain similar constraints on the eclipse depth.
@@
The eclipse model was generated using the package jaxoplanet (Hattori et al. 2024), which implements the equations of Mandel & Agol (2002) in JAX (Bradbury et al. 2018). The eclipse model was fixed for all simulations with the mean values of , , , from literature values of LHS 1140c (as listed in Table 10). The central eclipse time was fixed to the midpoint of the time series and the eccentricity was fixed to zero. For all fits of the simulated data, the eclipse depth was the only eclipse model parameter being fit for with every other eclipse parameter held fixed to the true values.
@@ 3.2 Simulating white noise and systematics
White noise was generated with the amplitude for each pixel determined by taking the standard deviation of each pixel time series in the first LHS 1140c eclipse. This noise was then convolved with the IPC convolution kernel from Equation 11 using values of , and , before then being added to the simulated signal. Background noise was also introduced to simulate residual background scatter not removed by a background subtraction. Each integration had a constant value added to all pixels, drawn from DN/s, which (given the flux in the target signal) works out to adding scatter with amplitude of ppm for a 5px radius circular aperture. We also added scatter which was constant along each row and column but independent between each row and column. This was added by drawing from DN/s for the rows and DN/s for the columns. Our analysis of the real observations did not constrain the presence of these row or column systematics but it was a possibility we wanted to include which both aperture photometry and the pixel-fitting method should be able to account for.
@@
We simulated the flux-conserved systematics by taking a random draw from a GP with an exponential kernel and a length scale of 35 minutes and adding it to one pixel and subtracting it from a neighbouring pixel. This was performed for each pair of neighbouring pixels (i.e. 8 times per pixel) and the amplitude of the draw was scaled to be the geometric mean of the flux on each of the pair of pixels multiplied by the constant ppm. By adding and subtracting the same signal, it means that these systematics are flux-conserving with the only non-flux-conserving effects due to the finite area of the aperture (i.e. there is ‘flux movement’ across the fixed aperture boundary), although the amplitude of this effect was small. For a pixel surrounded by eight pixels with the same flux values, the amplitude of systematics in this pixel would work out to ppm. We note that was meant to be set to 700 ppm - consistent with the values recovered in the observations - but the addition and subtraction was accidentally performed twice for each pair of pixels, effectively increasing the amplitude of this effect by. We generated 200 of these simulations and we refer to them as the flux-conserved systematics (FCS) scenario. The other two sets of simulations included additional systematics which were not flux-conserved. The second set of simulations – called the independent pixel systematics (IPS) scenario – took each simulation from the FCS scenario and injected a ppm height scale systematic into a single pixel time series. The pixel was randomly chosen for each simulation to be one of the nine brightest pixels at the centre of the PSF. As these pixels tend to make up about of total flux, this results in a ppm systematic in the aperture extracted light curve. While aperture extraction with this pixel masked out may be a good approach, the flux-conserved systematics make it difficult to distinguish these additional systematics from flux-conserved systematics. This can be seen in Figure 3, which features a plot of the pixel light curves centred on the PSF for the first simulated dataset in the FCS scenario and the IPS scenario.
@@
The third set of simulations added a common systematic shape to all pixels at an amplitude equal to ppm multiplied by the flux on each pixel. Here, this is referred to as the common systematics (CS) scenario. This could represent real astrophysical systematics such as a flare or instrumental systematics which affect the throughput of the observation. The height scale was chosen as it resulted in a similar amplitude of systematics in an aperture extracted light curve as the IPS scenario, but with the systematics spread out over all pixels instead of originating entirely from a single pixel.
@@
For each scenario, the synthetic data were aperture extracted with a radius of 5 pixels centred on the best-fit centroid for each integration – as performed for the real observations. This radius minimised scatter in the aperture extracted light curves compared to radii of 4.5 or 5.5 pixels, matching the real observations. In Figure 4, we give an example of a simulated dataset from the FCS and IPS scenarios after aperture extraction. The pixel-fitting method only included pixels with centres within 5 pixels from the PSF centre, resulting in 81 pixels being included, which is similar to the 78.5px2 area for a circular aperture with a radius of 5 pixels.
@@ 3.3 Different fitting approaches tested
We used MCMC to recover the eclipse depth with each method, modelling the posterior as Gaussian and using the mean and standard deviation of the chains. For each method, we ran four chains of No U-Turn Sampling (NUTS; Hoffman & Gelman 2014) as implemented in PyMC (Salvatier et al. 2016) – with each chain having 1000 tuning steps and 1000 draws. We also used slice sampling (Neal 2003) in combination with NUTS, by performing blocked Gibbs sampling (Jensen et al. 1995) on some parameters close to their prior bounds (e.g. , , , , , ). This was performed because it was found that running NUTS alone could sometimes result in chains getting stuck on these parameters. The Gelman-Rubin statistic (Gelman & Rubin 1992) was used to measure convergence of the chains. A slightly higher than normal threshold of was used to determine convergence for each simulation in order to reduce computational costs arising from the large number of MCMCs performed.
For each set of simulations, we tested aperture extraction with and without a GP. These MCMCs fit for the eclipse depth, , the flux out-of-(secondary) transit, , the white noise amplitude, , and in the case of the GP fits also the height scale, , and length scale, , of the GP. We also tested using a GP with fixed to the true simulated value of 35 minutes, as this may be a fairer comparison to the pixel-fitting method – which obtains extra information about the time length scale as it fits for flux-conserved systematics with the same time length scale.
For all pixel-fitting approaches, we fit for the eclipse depth, , baseline flux for each pixel, (for 81 pixels), white noise amplitude for each pixel , change in centroid position at each integration and (for 137 integrations), IPC parameters , , and , background scatter parameters, , and , height scale and time length scale of flux-conserving systematics and . This describes the ‘only flux-conserving systematics’ approach (abbreviated as ‘only FCS’) and includes 445 free parameters which were all fit for. We also included a height scale for systematics with a common shape in all pixels for the ‘common systematics’ approach and height scales for independent systematics in each pixel for the ‘independent systematics’ approach. Finally, as this independent systematics approach turned out to give quite conservative predictions, we also reran each of these approaches where only the pixels recovered to have large height scales (specifically where the percentile value was greater than a threshold of ppm) were fit with independent height scales and the remaining pixels all shared a joint height scale . We note these pixel systematics were still treated as independent but with a shared amplitude (i.e. they could still have completely different shapes). This is listed in tables as the ‘shared independent systematics’ approach. We also included methods which could account for both common and independent pixel systematics as these were robust to a range of systematics.
@@ 3.4 Comparing the performance of each method
We measured the precision of each method using the root mean square error (RMSE) of the mean eclipse depth for each simulation. The smaller the RMSE, the more precise the eclipse depth measurement is. If the method is robust, then the average standard deviation of the eclipse depth measurement, , should approximately match the RMSE. If the standard deviation is too big then our results are overly conservative and if it is too small then the results are unreliable. This can be summarised with the reduced chi-squared statistic , which should equal unity when the uncertainties are correctly estimated, less than one for uncertainties which are too large/conservative and greater than one for uncertainties which are too small/unreliable333This assumes the posteriors for the eclipse depths are Gaussian. For each scenario (consisting of 200 simulations), it is calculated as follows:
Here, ppm is the true simulated eclipse depth and , are the mean and standard deviation of the recovered eclipse depth for the simulation.
@@ 3.5 Flux-conserved systematics (FCS) scenario results
This scenario only contained flux-conserved systematics, therefore, aperture extraction should result in very clean light curves and a GP should not be required to robustly analyse these observations. However, for real observations we can never be sure that the light curves are completely free from systematics and so we are interested in how well each method performs in this scenario. The methods that marginalise over systematics will naturally tend to give more conservative constraints on the eclipse depths (even when systematics are not present); this is true both for aperture extraction with a GP and for various pixel-fitting approaches. However, this will be balanced out by their more reliable performance when systematics are present.
The results in Table 1 demonstrate this very clearly. The most precise methods (i.e. smallest RMSE) are aperture extraction without a GP and the pixel-fitting approach which only accounts for flux-conserved systematics. They both exhibit values close to unity across these simulations, showing their uncertainty estimates are reliable. Aperture extraction with a GP has a slightly worse precision and larger uncertainties. The significantly smaller value of shows that this method is quite conservative. Pixel-fitting only accounting for common systematics performs similarly.
However, we see that the other pixel-fitting approaches which can account for independent systematics in each pixel have worse accuracy and even larger uncertainty estimates. This is similar to aperture extraction with a GP and in general we do expect a more flexible systematics model to result in larger errorbars as it marginalises over a wider space of models (Gibson 2014). The pixel-fitting method is particularly flexible as it can fit for potential systematics in every pixel. Despite the vast majority of these pixel systematics height scales being consistent with the minimum prior bound, they still marginalise over a range of amplitudes which are consistent with the data. See Figure 5 for the posterior of each independent pixel systematics height scale, , for the first simulation in the FCS and IPS scenarios.
We see that sharing the amplitude of the smallest independent pixel systematics height scales does reduce how conservative the method is, with smaller average uncertainties and a value closer to one. Accounting for common systematics as well as independent pixel systematics offers the most conservative result; however this method is the most flexible and can be robustly applied to all three sets of simulations generated.
@@ 3.6 Independent pixel systematics (IPS) scenario results
For aperture extraction with a GP, the height scale of systematics is not constrained to be greater than the minimum prior bound in many of these simulations. This means that varying the location of this prior bound will increase or decrease the weighting of the area of the posterior where systematics are negligible. For example, we get if we increase the minimum bound to and if we instead lower the minimum bound to , we get . Increasing the minimum prior bound on the height scale can be interpreted as a more conservative analysis where the presence of systematics is more likely, and this would help lower the here. However, setting also lowers the value for the first set of simulations to just , which was already overly conservative at . Overall, for our normal prior bound of , the from combining the FCS and IPS scenarios is when fitting for and when fixing it; thus, the method performs well over an even mix of clean and systematics-contaminated datasets.
The pixel-fitting methods which account for common systematics on top of accounting for independent pixel systematics have slightly lower precision and larger uncertainties. However, they are accounting for additional systematics that could potentially exist in real data – such as those from variations in throughput and detector settling – so this is a more conservative approach that might make sense when we do not know the origin of systematics in the data.
@@ 3.7 Common systematics (CS) scenario results
Finally, the simulations in the common systematics scenario differ from the IPS scenario in that the systematics were injected into all pixels with a common shape instead of into a single pixel. The amplitude of these systematics was scaled by the flux in each pixel, so this could represent systematics inherent from the signal, such as stellar granulation or flaring. It could also arise from factors affecting the throughput which vary over time. The amplitude of this was such that aperture extraction with a GP performs similarly to the IPS scenario but with a different systematics origin. In the IPS scenario, the pixel-fitting approaches could outperform aperture extraction by weighting towards the cleaner pixels, however, this will not work when all the pixels suffer from a similar proportion of systematics. Pixel-fitting should therefore perform similarly to aperture extraction with a GP.
@@ 3.8 Summary of the simulations
The results demonstrate that pixel-fitting can outperform aperture extraction when systematics are large in a particular pixel, but the method is overly conservative when only flux-conserving systematics are present. It also matches the performance of aperture extraction with a GP when systematics have a common shape across all pixels. See Figure 7 for a summary of the results in Tables 1, 2, and 3 for aperture extraction with and without a GP as well as for the general pixel-fitting approach of accounting for common systematics across all pixels and systematics in individual pixels. We can see that compared with aperture extraction with a GP, pixel-fitting performs very similarly when systematics are common between all pixels but much better when there are systematics present in specific pixels; for instance, after a cosmic ray hit (as shown in Figure 1).
@@ 4 Observations
The three eclipses of LHS 1140c were observed on November 2023, July 2024, and July 2024. Each eclipse used the SUB256 subarray, read 36 groups for each integration and had 1,262 integrations. The integrations had an 11.1s cadence, with each observation lasting 233 minutes. This followed a strategy of leaving 60 minutes of flexible start time plus 30 minutes of settling time followed by observing for two eclipse durations, with one eclipse duration ( minutes) before and after expected mid-eclipse for a circular orbit. There was also an additional 9 minutes of padding to help catch slightly more eccentric orbits.
@@ 4.1 Primary data reduction
Our reduction began with a download of the uncalibrated ‘uncal’ FITS files for our three eclipses from the Mikulski Archive for Space Telescopes (MAST), which had already been pre-processed by version 2024_1a of the Science Data Processing (SDP) system. We ran the JWST calibration pipeline version 1.15.1 with the default settings for steps 1 and 2, except we changed the jump detection threshold to 7, excluded both the first and last groups from each integration when reading up-the-ramp (the default is to only exclude the last group) and skipped the photom step in stage 2 to keep the flux units in data numbers per second (DN/s). We excluded the first and last groups to avoid a known issue where these groups can include difficult to correct reset effects (Morrison et al. 2023). The jump detection threshold was decided upon by calculating the average amplitude of white noise recovered across all three eclipse fits when only white noise and an exponential ramp + linear model was fit. It was found that 7 minimised this out of integer thresholds from .
One issue we encountered was that running the pipeline separately on each uncal data segment resulted in offsets in the flux time series between segments, despite these segments not representing any actual change in the observation but instead are simply a way of reducing the total filesize of any given file. To remedy this, for all our reductions we first combined the uncal segments together into a single FITS file and then ran the JWST pipeline on this combined uncal file, which avoided these offsets.